# Teaching Code LLMs to Use Autocompletion Tools in Repository-Level Code Generation

### v2

@@ Abstract
Recent code large language models (LLMs) have shown promising performance in generating standalone functions. However, they face limitations in repository-level code generation due to their lack of awareness of repository-level dependencies (e.g., user-defined attributes), resulting in dependency errors such as undefined-variable and no-member errors. In this work, we introduce ToolGen, an approach that integrates autocompletion tools into the code LLM generation process to address these dependencies. ToolGen comprises two main phases: Trigger Insertion and Model Fine-tuning (Offline), and Tool-integrated Code Generation (Online). During the offline phase, ToolGen augments functions within a given code corpus with a special mark token, indicating positions to trigger autocompletion tools. These augmented functions, along with their corresponding docstrings, are then used to fine-tune a selected code LLM. In the online phase, ToolGen iteratively generates functions by predicting tokens step-by-step using the fine-tuned LLM. Whenever a mark token is encountered, ToolGen invokes the autocompletion tool to suggest code completions and selects the most appropriate one through constrained greedy search.
We conduct comprehensive experiments to evaluate ToolGen’s effectiveness in repository-level code generation. To facilitate this evaluation, we create a benchmark comprising 680 real-world code repositories and introduce two new repository-level metrics: Dependency Coverage and Static Validity Rate. The results demonstrate that ToolGen significantly improves Dependency Coverage by 15.2% to 45.8% and Static Validity Rate by 10.9% to 42.2% across three distinct code LLMs, while maintaining competitive performance in widely-recognized similarity metrics such as BLEU-4, CodeBLEU, and Edit Similarity. Furthermore, our generalizability evaluation confirms ToolGen’s consistent performance when applied to diverse code LLMs, including various model architectures and scales.
@@ I Introduction
Code generation has been a longstanding focal point in the field of software engineering. Recent advancements have introduced a variety of code large language models (LLMs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] constructed upon the Transformer model architecture [14]. These models are either pre-trained or fine-tuned on extensive code corpora, enabling them to automatically generate code based on provided natural language descriptions. These code LLMs have demonstrated notable effectiveness in the generation of code blocks or functions. For instance, CodeLlama [12], built upon the foundational Llama2 model [15], has achieved state-of-the-art results among open code LLMs (e.g., CodeGen [8] and StarCoder [11]), on benchmarks like HumanEval [3] and MBPP [16] that focus on standalone functions.
@@
In fact, recent research has delved into the integration of external tools into the generation process of LLMs to mitigate their limitations in constrained generation scenarios. One noteworthy example is ToolFormer [20], which creates an augmented dataset to instruct LLMs on invoking existing arithmetic calculators. This integration effectively reduces errors in generated text involving arithmetic calculations. Building upon ToolFormer’s inspiration, Zhang et al. [21] introduce ToolCoder, an approach designed to teach LLMs how to utilize information-retrieval-based (IR-based) API search tools during the code generation process. While ToolCoder targets the generation of functionally correct standalone functions and demonstrates promising results, the integrated IR-based API search tools do not consider repository-level dependencies, limiting their potential in resolving dependency errors. Another relevant example of harnessing external tools is Repilot [22], which leverages code completion tools to filter out impractical suggestions made by LLMs in the context of automatic program repairing (APR). Unlike repository-level code generation, Repilot’s primary focus is on generating valid bug-fix patches rather than entire functions.
@@
To tackle the challenges, we propose ToolGen, an approach to integrate autocompletion tools into the generation process of code LLMs to support repository-level code generation. ToolGen has two main phases: Trigger Insertion and Model Fine-tuning (Offline), and Tool-integrated Code Generation (Online). In the offline phase, ToolGen analyzes source files within a corpus of code repositories, creating abstract syntax trees (ASTs) and extracting function definitions. It augments these functions by inserting a special token, <COMP>, signifying the positions to trigger autocompletion tools. The insertion positions are established by navigating through the functions and identifying the identifiers that can be recommended by autocomplete tools. These augmented functions, paired with their respective docstrings, are then employed to fine-tune a selected code LLM. In the online phase, ToolGen iteratively constructs a function based on a provided docstring by predicting tokens step-by-step through the fine-tuned LLM. Whenever a <COMP> token is encountered, ToolGen invokes the autocompletion tool to suggest code completions, drawing from the current repository context. Subsequently, it identifies the most appropriate suggestion through a constraint greedy search algorithm, appending this selected suggestion to the current tokens. This process continues as it predicts tokens until a specified termination condition is satisfied.
We conduct extensive experiments to evaluate the effectiveness of ToolGen in repository-level code generation. To facilitate this evaluation, we first construct a benchmark, which includes 12,744 Python functions from 680 real-world code repositories and defines two new repository-level metrics, namely Dependency Coverage and Static Validity Rate. Dependency Coverage quantifies the proportion of repository-level dependencies present in ground-truth functions and successfully covered by the generated functions, while Static Validity Rate measures the percentage of generated functions that pass a dependency error check. The evaluation results demonstrate that ToolGen exhibits comparable performance in widely-recognized similarity metrics such as BLEU-4, CodeBLEU, and Edit Similarity. Importantly, ToolGen achieves significant improvements in Dependency Coverage, ranging from 15.2% to 45.8%, and Static Validity Rate, spanning from 10.9% to 42.2%, across three distinct code LLMs. Moreover, the results from our generalizability evaluation confirm that ToolGen consistently performs well across a variety of code LLMs, with different model architectures and scales.
@@
  • ToolGen, an novel approach that seamlessly integrates autocompletion tools into the generation process of code LLMs, thereby enhancing repository-level code generation.
  • An Augmented Dataset, which comprises 256,870 Python functions sourced from a diverse selection of 12,231 code repositories. Each function is augmented with a special token, <COMP>, which signifies positions suitable for invoking autocompletion tools.
  • An Evaluation Benchmark, which encompasses 12,744 Python functions drawn from 680 real-world code repositories, along with the introduction of two novel repository-level metrics: Dependency Coverage and Static Validity Rate.
  • Extensive Experimental Results, which affirm the efficacy of ToolGen in repository-level code generation. ToolGen demonstrates substantial improvements in Dependency Coverage, ranging from 15.2% to 45.8%, and Static Validity Rate, spanning from 10.9% to 42.2%, across three distinct code LLMs.
@@ II-A Code LLMs
Typically, there are two main categories of code LLMs that can be employed for code generation. These categories include decoder-only models and encoder-decoder models, each of which conducts the code generation process base on a given docstring as outlined below:
  • Decoder-only Models: Illustrated in Figure 1(a), decoder-only code LLMs, such as CodeGPT [23] and CodeLlama [12], consist solely of a decoder component derived from the Transformer architecture [14]. An employed decoder-only model first tokenizes the input docstring into a sequence of tokens. Subsequently, it feeds this token sequence into the model’s decoder and proceeds to predict a function token-by-token, based on the context provided by the docstring and previously predicted tokens.
  • Encoder-Decoder Models: As depicted in Figure 1(b), encoder-decoder code LLMs, such as CodeT5 [4] and CodeT5+ [5], encompass both the encoder and decoder components of the Transformer architecture. In this case, the employed model also tokenizes the docstring into a token sequence, but the sequence is first processed by the model’s encoder. The model’s decoder is then tasked with predicting a function token-by-token, relying on the representation produced by the encoder and the context provided by the preceding tokens.
@@
Here, $\Sigma_{char}^{*}$ represents a character sequence of either a docstring or a code snippet, and $\mathbb{V}^{*}$ corresponds to the resulting sequence of tokens drawn from $\mathbb{V}$.
@@
In this context, the two input token sequences ($\mathbb{V}^{*}$) represent a docstring and an incomplete function, respectively, while $[0,1]^{|\mathbb{V}|}$ signifies a probability distribution encompassing $|\mathbb{V}|$ probabilities $[0,1]$. Here, $|\mathbb{V}|$ is the size (token numbers) of the vocabulary $\mathbb{V}$.
@@ III-A Overview
In trigger insertion and model fine-tuning, ToolGen parses each source file in the given code repositories into an abstract syntax tree (AST) and then extract extracts function definitions from the AST; For each extract function definition, ToolGen then utilizes an autocompletion tool to augment it with the special token <COMP> to mark the positions to invoke the tool, and then assembles a pair of docstring and augmented function; After process all code repositories, ToolGen employs the resulting pairs of docstrings and augmented functions to fine-tune a code LLM, resulting in a fine-tuned code LLM that can predict <COMP> at suitable positions to trigger the autocompletion tool.
In tool-integrated code generation, ToolGen generates a token sequence to form a function by an iterative process, in which, at each step, one or multiple tokens are yielded by the fine-tuned code LLM and the employed autocompletion tool. At certain step, ❶ the fine-tuned code LLM takes the given docstring and the incomplete function as inputs and predicts the next token; The predicted token is appended to the incomplete function; ❷ If the predicted token equals <COMP>, the autocomplete tool is triggered and a list of completion suggestions is returned based on the current repository context; ❸ ToolGen then selects the most suitable one from the suggestions with the fine-tuned code LLM and appends the selected suggestion to the incomplete function.
@@ III-B1 Trigger Insertion
$\mathcal{F}_{aug}\leftarrow\textsc{getSignature}(node)$   // signature
@@ 4             $\mathbb{C}\leftarrow\textsc{tool-complete}(\mathcal{R},\mathcal{P})$
Next, we assemble a tuple $(\mathcal{D},\mathcal{F}_{aug})$, in which $\mathcal{D}$ corresponds to the docstring of the parsed function. Note that functions lacking corresponding docstrings are omitted from our process as our repository-level code generation relies on textual descriptions as input. Once we complete the processing of all code repositories, we accumulate an augmented dataset that contains a substantial number of these data tuples.
@@ III-B2 Model Fine-tuning
During the fine-tuning process, we supply the collected docstrings and augmented functions to optimize the parameters of the employ code LLM (base model), adhering to established practices in code generation tasks. Specifically, for each pair consisting of a docstring $\mathcal{D}$ and an augmented function $\mathcal{F}_{aug}$, both are tokenized into sequences of tokens and subsequently fed into the base model to undergo the token-by-token generation process described in Section II-A. At each step, a cross-entropy loss is computed between the predicted probability distribution of the next token and the ground-truth next token present in $\mathcal{F}_{aug}$.
@@ III-C1 Overall Process
Algorithm 2 outlines the overall tool-integrated generation process, comprising three crucial parts based on the fine-tuned code LLM and the employed autocompletion tool: ❶ Next Token Prediction, ❷ Code Autocompletion, and ❸ Suggestion Selection. This algorithm takes a code repository $\mathcal{R}$, an insertion position $\mathcal{P}$, and a docstring $\mathcal{D}$ as inputs and follows an iterative process to generate a token sequence, ultimately constructing a function denoted as $\mathcal{F}$. Here, the tokens are drawn from the expanded vocabulary $\mathbb{V}$ defined in Equation 1.
@@
Input: Repository $\mathcal{R}$, Docstring $\mathcal{D}$, Insertion Position $\mathcal{P}$
@@ 1 Procedure llm-select($\mathcal{D},\mathcal{F}$, $\mathbb{C}$):
       $trie\leftarrow Trie()$   // prefix tree
@@ IV Evaluation Setup
To evalute the effectiveness of ToolGen in repository-level code generation, we conduct a comprehensive set of experiments.
@@ IV-A Research Questions
  • RQ1 - Similarity Metrics: How closely does the code generated by ToolGen align with the ground truth when assessed using common similarity metrics?
  • RQ2 - Repository-level Metrics:
    • RQ2.a: What extent does the generated code cover repository-level dependencies, including user-defined functions and attributes?
    • RQ2.b: To what degree can ToolGen reduce errors, such as no-member and undefined-variable errors, arising from repository-level dependencies?
  • RQ3 - Generalizability: Is ToolGen effective in code generation when applied to different code LLMs?

### v3

@@ Abstract
Recent code large language models (LLMs) have shown promising performance in generating standalone functions. However, they face limitations in repository-level code generation due to their lack of awareness of repository-level dependencies (e.g., user-defined attributes), resulting in dependency errors such as undefined-variable and no-member errors. In this work, we introduce ToolGen, an approach that integrates autocompletion tools into the code LLM generation process to address these dependencies. ToolGen comprises two main phases: Trigger Insertion and Model Fine-tuning (Offline), and Tool-integrated Code Generation (Online). During the offline phase, ToolGen augments functions within a given code corpus with a special mark token, indicating positions to trigger autocompletion tools. These augmented functions, along with their corresponding descriptions, are then used to fine-tune a selected code LLM. In the online phase, ToolGen iteratively generates functions by predicting tokens step-by-step using the fine-tuned LLM. Whenever a mark token is encountered, ToolGen invokes the autocompletion tool to suggest code completions and selects the most appropriate one through constrained greedy search.
We conduct comprehensive experiments to evaluate ToolGen’s effectiveness in repository-level code generation across three distinct code LLMs: CodeGPT, CodeT5, and CodeLlama. To facilitate this evaluation, we create a benchmark comprising 671 real-world code repositories and introduce two new dependency-based metrics: Dependency Coverage and Static Validity Rate. The results demonstrate that ToolGen significantly improves Dependency Coverage by 31.4% to 39.1% and Static Validity Rate by 44.9% to 57.7% across the three LLMs, while maintaining competitive or improved performance in widely recognized similarity metrics such as BLEU-4, CodeBLEU, Edit Similarity, and Exact Match. On the CoderEval dataset, ToolGen achieves improvements of 40.0% and 25.0% in test pass rate (Pass@1) for CodeT5 and CodeLlama, respectively, while maintaining the same pass rate for CodeGPT. ToolGen also demonstrates high efficiency in repository-level code generation, with latency ranging from 0.63 to 2.34 seconds for generating each function. Furthermore, our generalizability evaluation confirms ToolGen’s consistent performance when applied to diverse code LLMs, encompassing various model architectures and scales.
@@ I Introduction
Code generation has been a longstanding focal point in the field of software engineering. Recent advancements have introduced a variety of code large language models (LLMs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] constructed upon the Transformer model architecture [14], achieving promising performance in code-related applications [15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. These models are either pre-trained or fine-tuned on extensive code corpora, enabling them to automatically generate code based on provided natural language descriptions. These code LLMs have demonstrated notable effectiveness in the generation of code blocks or functions. For instance, CodeLlama [12], built upon the foundational Llama2 model [25], has achieved state-of-the-art results among open code LLMs (e.g., CodeGen [8] and StarCoder [11]), on benchmarks like HumanEval [3] and MBPP [26] that focus on standalone functions.
@@
In fact, recent research has delved into the integration of external tools into the generation process of LLMs to mitigate their limitations in constrained generation scenarios. One noteworthy example is ToolFormer [30], which creates an augmented dataset to instruct LLMs on invoking existing arithmetic calculators. This integration effectively reduces errors in generated text involving arithmetic calculations. Building upon ToolFormer’s inspiration, Zhang et al. [31] introduce ToolCoder, an approach designed to teach LLMs how to utilize information-retrieval-based (IR-based) API search tools during the code generation process. While ToolCoder targets the generation of functionally correct standalone functions and demonstrates promising results, the integrated IR-based API search tools do not consider repository-level dependencies, limiting their potential in resolving dependency errors. Additionally, ToolFormer and ToolCoder are unable to handle scenarios where the tools return multiple candidates. Another relevant example of harnessing external tools is Repilot [32], which leverages code completion tools to filter out impractical suggestions made by LLMs in the context of automatic program repairing (APR). Unlike repository-level code generation, Repilot’s primary focus is on generating valid single-hunk bug-fix patches rather than entire functions. When applying Repilot to function-level code generation, the autocompletion tools are frequently triggered unnecessarily, resulting in significant overhead and impracticality.
@@
To tackle the challenges, we propose ToolGen, an approach to integrate autocompletion tools into the generation process of code LLMs to support repository-level code generation. ToolGen has two main phases: Trigger Insertion and Model Fine-tuning (Offline), and Tool-integrated Code Generation (Online). In the offline phase, ToolGen analyzes source files within a corpus of code repositories, creating abstract syntax trees (ASTs) and extracting function definitions. It augments these functions by inserting a special token, <COMP>, signifying the positions to trigger autocompletion tools. The insertion positions are established by navigating through the functions and identifying the identifiers that can be recommended by autocomplete tools. These augmented functions, paired with their respective descriptions, are then employed to fine-tune a selected code LLM. In the online phase, ToolGen iteratively constructs a function based on a provided description by predicting tokens step-by-step through the fine-tuned LLM. Whenever a <COMP> token is encountered, ToolGen invokes the autocompletion tool to suggest code completions, drawing from the current repository context. Subsequently, it identifies the most appropriate suggestion through a constraint greedy search algorithm, appending this selected suggestion to the current tokens. This process continues as it predicts tokens until a specified termination condition is satisfied.
We conduct extensive experiments to evaluate the effectiveness of ToolGen in repository-level code generation across three distinct code LLMs, namely, CodeGPT [33], CodeT5 [4], and CodeLlama [12]. To facilitate this evaluation, we first construct a benchmark, which includes 12,406 Python functions from 671 real-world code repositories and 176 coding tasks from CoderEval dataset [27]. We define two new repository-level metrics, namely Dependency Coverage and Static Validity Rate. Dependency Coverage quantifies the proportion of repository-level dependencies present in ground-truth functions and successfully covered by the generated functions, while Static Validity Rate measures the percentage of generated functions that pass a dependency error check. The evaluation results on the 12,406 functions demonstrate that ToolGen exhibits comparable or improved performance in widely-recognized similarity metrics such as BLEU-4, CodeBLEU, Edit Similarity, and Exact Match. Importantly, ToolGen achieves significant improvements in Dependency Coverage, ranging from 31.4% to 39.1%, and Static Validity Rate, spanning from 44.9% to 57.7%, across the three code LLMs. On the 176 tasks derived from CoderEval, ToolGen achieves improvements of 40.0% and 25.0% in test pass rate (Pass@1) for CodeT5 and CodeLlama, respectively, while maintaining the same pass rate for CodeGPT. ToolGen also demonstrates high efficiency in repository-level code generation, with average latency ranging from 0.63 to 2.34 seconds, attributed to offline fine-tuning with trigger insertion. Moreover, the results from our generalizability evaluation confirm that ToolGen consistently performs well across a variety of code LLMs, with different model architectures and scales.
@@
  • ToolGen, an novel approach that seamlessly integrates autocompletion tools into the generation process of code LLMs, which consists of Trigger Insertion and Model Fine-tuning (Offline), and Tool-integrated Code Generation (Online). ToolGen seamlessly integrates the autocompletion tool into the generation process of code LLMs, thereby enhancing repository-level code generation. The offline phase results in an Augmented Dataset, which comprises 249,298 Python functions sourced from a diverse selection of 12,231 code repositories. Each function is augmented with a special token, <COMP>, which signifies positions suitable for invoking autocompletion tools.
  • An Evaluation Benchmark, which encompasses 12,406 Python functions drawn from 671 real-world code repositories and 176 coding tasks with test cases derived from CoderEval, along with the introduction of two novel repository-level metrics: Dependency Coverage and Static Validity Rate.
  • Extensive Experimental Results, which affirm the efficacy of ToolGen in repository-level code generation. ToolGen demonstrates substantial improvements in Dependency Coverage, ranging from 31.4% to 39.1%, and Static Validity Rate, spanning from 44.9% to 57.7%, across three distinct code LLMs. Additionally, ToolGen achieves 40% and 25% improvements in test pass rate for CodeT5 and CodeLlama, respectively, with high generation efficiency.
@@ II-A Code LLMs
Typically, there are two main categories of code LLMs that can be employed for code generation. These categories include decoder-only models and encoder-decoder models, each of which conducts the code generation process base on a given description as outlined below:
  • Decoder-only Models: Illustrated in Figure 2(a), decoder-only code LLMs, such as CodeGPT [33] and CodeLlama [12], consist solely of a decoder component derived from the Transformer architecture [14]. An employed decoder-only model first tokenizes the input description into a sequence of tokens. Subsequently, it feeds this token sequence into the model’s decoder and proceeds to predict a function token-by-token, based on the context provided by the description and previously predicted tokens.
  • Encoder-Decoder Models: As depicted in Figure 2(b), encoder-decoder code LLMs, such as CodeT5 [4] and CodeT5+ [5], encompass both the encoder and decoder components of the Transformer architecture. In this case, the employed model also tokenizes the description into a token sequence, but the sequence is first processed by the model’s encoder. The model’s decoder is then tasked with predicting a function token-by-token, relying on the representation produced by the encoder and the context provided by the preceding tokens.
@@
Here, $\Sigma_{char}^{*}$ represents a character sequence of either a description or a code snippet, and $\mathbb{V}^{*}$ corresponds to the resulting sequence of tokens drawn from $\mathbb{V}$.
@@
In this context, the two input token sequences ($\mathbb{V}^{*}$) represent a description and an incomplete function, respectively, while $[0,1]^{|\mathbb{V}|}$ signifies a probability distribution encompassing $|\mathbb{V}|$ probabilities $[0,1]$. Here, $|\mathbb{V}|$ is the size (token numbers) of the vocabulary $\mathbb{V}$.
@@ III-A Overview
In trigger insertion and model fine-tuning, ToolGen parses each source file in the given code repositories into an abstract syntax tree (AST) and then extracts function definitions from the AST; For each extract function definition, ToolGen then utilizes an autocompletion tool to augment it with the special token <COMP> to mark the positions to invoke the tool, and then assembles a pair of description and augmented function; After process all code repositories, ToolGen employs the resulting pairs of descriptions and augmented functions to fine-tune a code LLM, resulting in a fine-tuned code LLM that can predict <COMP> at suitable positions to trigger the autocompletion tool.
In tool-integrated code generation, ToolGen generates a token sequence to form a function by an iterative process, in which, at each step, one or multiple tokens are yielded by the fine-tuned code LLM and the employed autocompletion tool. At certain step, ❶ the fine-tuned code LLM takes the given description and the incomplete function as inputs and predicts the next token; The predicted token is appended to the incomplete function; ❷ If the predicted token equals <COMP>, the autocomplete tool is triggered and a list of completion suggestions is returned based on the current repository context; ❸ ToolGen then selects the most suitable one from the suggestions with the fine-tuned code LLM and appends the selected suggestion to the incomplete function.
@@ III-B1 Trigger Insertion
$\mathcal{F}_{aug}\leftarrow\textsc{getSignature}(node)$
  // signature
@@ 4             $\mathbb{C}\leftarrow\textsc{tool-complete}(\mathcal{R},\mathcal{P})$
Next, we assemble a tuple $(\mathcal{D},\mathcal{F}_{aug})$, in which $\mathcal{D}$ corresponds to the concatenation of the signature and docstring of the parsed function. Note that functions lacking corresponding docstrings are omitted from our process as our repository-level code generation relies on textual descriptions as input. Once we complete the processing of all code repositories, we accumulate an augmented dataset that contains a substantial number of these data tuples.
Note that our trigger insertion method can be applied to arbitrary code and is not limited to function bodies alone. Currently, we focus exclusively on function bodies, as our primary application scenario involves generating code based on the given natural language descriptions. Extracting descriptions for code blocks outside functions for model training and evaluation is challenging, due to the difficulty in determining the scope of line comments [34, 35]. Therefore, we solely consider function bodies, where corresponding descriptions can be readily obtained from function docstrings.
@@ III-B2 Model Fine-tuning
During the fine-tuning process, we supply the collected descriptions and augmented functions to optimize the parameters of the employ code LLM (base model), adhering to established practices in code generation tasks. Specifically, for each pair consisting of a description $\mathcal{D}$ and an augmented function $\mathcal{F}_{aug}$, both are tokenized into sequences of tokens and subsequently fed into the base model to undergo the token-by-token generation process described in Section II-A. At each step, a cross-entropy loss is computed between the predicted probability distribution of the next token and the ground-truth next token present in $\mathcal{F}_{aug}$.
@@ III-C1 Overall Process
Algorithm 2 outlines the overall tool-integrated generation process, comprising three crucial parts based on the fine-tuned code LLM and the employed autocompletion tool: ❶ Next Token Prediction, ❷ Code Autocompletion, and ❸ Suggestion Selection. This algorithm takes a code repository $\mathcal{R}$, an insertion position $\mathcal{P}$, and a description $\mathcal{D}$ as inputs and follows an iterative process to generate a token sequence, ultimately constructing a function denoted as $\mathcal{F}$. Here, the tokens are drawn from the expanded vocabulary $\mathbb{V}$ defined in Equation 1.
@@
Input: Repository $\mathcal{R}$, Description $\mathcal{D}$, Insertion Position $\mathcal{P}$
@@ 1 Procedure llm-select($\mathcal{D},\mathcal{F}$, $\mathbb{C}$):
       $trie\leftarrow Trie()$
        // prefix tree
@@ IV Evaluation Setup
To evaluate the effectiveness and efficiency of ToolGen in repository-level code generation, we conduct a comprehensive set of experiments.
@@ IV-A Research Questions
  • RQ1 - Similarity-based Effectiveness: How closely does the code generated by ToolGen align with the ground truth when assessed using common similarity metrics?
  • RQ2 - Dependency-based Effectiveness: To what degree can ToolGen cover repository-level dependencies and reduce dependency errors, including those related to user-defined functions and attributes?
  • RQ3 - Execution-based Effectiveness: How effectively can ToolGen generate functionally correct functions that pass test cases?
  • RQ4 - Efficiency: What is the average time ToolGen takes to generate functions?
  • RQ5 - Generalizability: Is ToolGen effective in code generation when applied to different code LLMs?